OpenStudy (anonymous):

domain:-1 0 1 2 3 4 range: -2 0 0 1 2 3 what is the rule?

OpenStudy (anonymous):

what does the domain differ from the range?

OpenStudy (anonymous):

domain is the x and range is the y'

OpenStudy (amistre64):

a rather lengthy way is to simply setup a 5th degree poly that hits all those points

OpenStudy (amistre64):

and i think i mixed up some x and y parts in that so its prolly not the correct stuff; but the general idea is sound :)

OpenStudy (amistre64):

-1 0 1 2 3 4 (x+1) (x-0) (x-1) (x-2) (x-3) (x-4) +a(x-0)(x-1)(x-2)(x-3)(x-4) +b(x+1)(x-1)(x-2)(x-3)(x-4) +c(x+1)(x-0)(x-2)(x-3)(x-4) +d(x+1)(x-0)(x-1)(x-3)(x-4) +e(x+1)(x-0)(x-1)(x-2)(x-4) +f(x+1)(x-0)(x-1)(x-2)(x-3) that might prove more exacting

OpenStudy (amistre64):

when x=-1, y=-2; everything zeros out but the a term, giving us \[-2=a(-1-0)(-1-1)(-1-2)(-1-3)(-1-4)\] \[\frac{-2}{(-1)(-2)(-3)(-4)(-5)}=a\] \[\frac{1}{60}=a\]

OpenStudy (amistre64):

I believe newton had a similar method for hitting all the points; just as lengthy tho setup so that each term zeros out to build up the coeffs as you go y = a + b(x+1) + c(x+1)(x-0) + d(x+1)(x-0)(x-1) +e(x+1)(x-0)(x-1)(x-2) + f(x+1)(x-0)(x-1)(x-2)(x-3) now, when x=-1, y=-2 such that -2 = a y = -2 when x=0, y=0 0 = -2 + b(1) ; b=2 y = -2 + 2x+2 + c(x+1)(x-0) + d(x+1)(x-0)(x-1) +e(x+1)(x-0)(x-1)(x-2) + f(x+1)(x-0)(x-1)(x-2)(x-3) y = 2x + c(x+1)(x-0) + d(x+1)(x-0)(x-1) +e(x+1)(x-0)(x-1)(x-2) + f(x+1)(x-0)(x-1)(x-2)(x-3) etc ...

OpenStudy (amistre64):

(1,0) 0 = 2 + 2c ; c = -1 y = -x^2 + x + d(x+1)(x-0)(x-1) +e(x+1)(x-0)(x-1)(x-2) + f(x+1)(x-0)(x-1)(x-2)(x-3) (2,1) 1 = -4 + 2 + 6d 3 = 6d ; d=2 y = 2x^3 - x^2 + 3x + e(x+1)(x-0)(x-1)(x-2) + f (x+1)(x-0)(x-1)(x-2)(x-3) etc

OpenStudy (amistre64):

err, d=1/2 :)